OpenStudy (lgbasallote):

Find: \[\huge \frac d{dx} [x \tan^{-1} (e^x)]\]

hartnn (hartnn):

didn't try product rule first then chain rule ?

OpenStudy (lgbasallote):

product rule first then chain rule?

hartnn (hartnn):

can u only diff. tan^-1 (e^x) ? or that part u need ?

hartnn (hartnn):

let y= tan^-1 (e^x) d/dx(xy) = xy' + yx' = xy' + y this product rule i was talking about....

OpenStudy (lgbasallote):

i know the product rule......but how to derive tan^-1 (e^x)

hartnn (hartnn):

know the derivative of tan^-1 x ?

OpenStudy (zugzwang):

Just differentiate tan^-1 x... implicitly?

OpenStudy (lgbasallote):

no idea

OpenStudy (zugzwang):

\[y = \tan^{-1}x\] \[x = \tan y\] Differentiate both sides with respect to x.

OpenStudy (lgbasallote):

nevermind...found it....seems using laplace transformation can work...

hartnn (hartnn):

d/dx (tan^-1 x) = 1/ (1+x^2)

OpenStudy (zugzwang):

And you once told me I was making things more complicated than they should be...

hartnn (hartnn):

O.o Laplace!

OpenStudy (lgbasallote):

that's why i made it simple @zugzwang

OpenStudy (zugzwang):

@hartnn already has the derivative... \[x = \tan y\]\[1 = \sec^{2}y \frac{dy}{dx}\]\[\frac{dy}{dx}=\frac{1}{\sec^{2}y}\]\[\frac{dy}{dx}=\frac{1}{\tan^{2}y+1}\] since x = tan y then \[\frac{dy}{dx}=\frac{1}{x^{2}+1}\]